(3.3.15) Based on sample​ data, newborn males have weights with a mean of 3243.4 g and a standard deviation of 574.9 g. Newborn females have weights with a mean of 3025.6 g and a standard deviation of 724.6 g. Who has the weight that is more extreme relative to the group from which they​ came: a male who weighs 1700g or a female who weighs 1700 ​g?

Since the z score for the male is z=? and the z score for the female is z=?​, the female/male has the weight that is more extreme.
1
Expert's answer
2020-11-23T18:29:21-0500

Z for male is "\frac{1700-3243.4}{574.9}=-2.68"

Z for female is "\frac{1700-3025.6}{724.6}=-1.83"

Since the z score for the male is z=-2.68 and the z score for the female is z=-1.83, the male has the weight that is more extreme.